An engine operates at a constant temperature of 110 degrees celsius through a reversible process. The engine work output is 5.30kJ and the heat loss is 6.20 kJ. Determine the change in entropy during the process.


1
Expert's answer
2021-07-06T01:16:29-0400

"\Delta S= \frac{Q_{out}}{T_{reservior}}= \frac{6200}{110+273}= 16.19 J/K"